OpenStudy (anonymous):

If BY = 6, YC = 10, AX = 18, then XC =

OpenStudy (anonymous):

heres a photo

Parth (parthkohli):

Concept of similar triangles is followed.

OpenStudy (anonymous):

I think the ratios of the sides we have to be taken...

Parth (parthkohli):

\(10:18 = 18:x\)

Parth (parthkohli):

Yes, @waterineyes

OpenStudy (anonymous):

\[\frac{BY}{AX} = \frac{YC}{XC}\] Or: \[\frac{6}{18}= \frac{10}{XC}\]

OpenStudy (anonymous):

Find \(XC\) from here..